OpenStudy (anonymous):

Solve -4x - 5 > -9 and describe the graph of the solution. x > 1; closed circle on 1, shading to the right x < 1; closed circle on 1, shading to the left x < 1; open circle on 1, shading to the left x > 1; open circle on 1, shading to the right

OpenStudy (anonymous):

dude. quit spamming please.

OpenStudy (anonymous):

-4x-5>-9 -4x>-4 x<1 when dividing by -4, you must flip the sign since its only < we can assume it is open circle now test x=0 -4(0)-5>-9 -5>-9 true x=1 -4(1)-5>-9 -9>-9 false open circle x=2 -4(2)-5>-9 -13>-9 false therefore shading to left